OpenStudy (richley):

Need help with Honors Pre-Calculus Questions. They have to do with the binomial theorem, DeMoivre's theorem, and a+ib form.

OpenStudy (richley):

1) ALREADY SOLVED, BUT NEED THIS ANSWER FOR THE OTHER 2 Use the binomial theorem to express (a+ib)^n in form of A+iB where A and B are summations. Answer: \[\sum_{j=0}^{\frac{ n-1 }{ 2 }}\left(\begin{matrix}n \\ 2j\end{matrix}\right)a^{n-2j}(-1)^jb ^{2j}+i \sum_{j=0}^{\frac{ n-1 }{ 2 }}a ^{n-(2j+1)}(-1)^jb ^{2j+1} \]

OpenStudy (richley):

2) Given \[t \in R, n \in N,\] use DeMoivre's theorem and the result of #1 to express: cis(nt) = cos(nt) + isin(nt) to where cos(nt) and sin(nt) are summations of powers of cos(t) and sin(t). Also express cos(nt) in powers of cos(t) only (using \[\cos^2 = 1-\sin^2\])

OpenStudy (richley):

3) Given \[t \in R\] Use the result of problem 2 to express, in simplest form: A) cos(5t) in terms of powers of cos(t) B) sin(5t) in terms of powers of sin(t)

OpenStudy (richley):

Thanks a ton for the help guys. These two problems are really confusing for me and I frankly don't know where to start.

OpenStudy (richley):

@ganeshie8

OpenStudy (richley):

Hey, sorry to tag you but you helped me a TON last time.

ganeshie8 (ganeshie8):

cis(nt) = cos(nt) + isin(nt) = (cost + isint)^n use binomial theorem ?

ganeshie8 (ganeshie8):

(cost + isint)^n = ?

ganeshie8 (ganeshie8):

You may use the result from part 1

OpenStudy (richley):

Yeah that's where demoivre's theory comes in right?

ganeshie8 (ganeshie8):

Notice, we're using the euler formula : \(e^{it} = \cos(t) + i\sin(t)\)

OpenStudy (richley):

means we can express it in terms of a power to n

ganeshie8 (ganeshie8):

Demoivre's theorem tells us \((e^{it})^n = (\cos(t) + i\sin(t))^n = \cos(nt) + i\sin(nt)\)

OpenStudy (richley):

Yeah

ganeshie8 (ganeshie8):

You're right, simply use the result from part 1 and compare the real part both sides

OpenStudy (richley):

Ah. So I just plugit in to the previous formula?

OpenStudy (richley):

So cos(t) would be "a" and sin(t) would be "b" and the n remains "n"

ganeshie8 (ganeshie8):

Yes : \[\cos(nt) + i\sin(nt) \\~\\ = (\cos t+ i\sin t)^n\\~\\ = \sum_{j=0}^{\frac{ n-1 }{ 2 }}\left(\begin{matrix}n \\ 2j\end{matrix}\right)(\cos t)^{n-2j}(-1)^j (\sin t) ^{2j}+i \sum_{j=0}^{\frac{ n-1 }{ 2 }}(\cos t)^{n-(2j+1)}(-1)^j)(\sin t) ^{2j+1}\]

ganeshie8 (ganeshie8):

Compare the real parts from first line and last lines

OpenStudy (richley):

Its just the +1 after the 2j, right?

ganeshie8 (ganeshie8):

The red parts are equal : \[\color{red}{\cos(nt)} + i\sin(nt) \\~\\ = (\cos t+ i\sin t)^n\\~\\ = \color{Red}{\sum_{j=0}^{\frac{ n-1 }{ 2 }}\left(\begin{matrix}n \\ 2j\end{matrix}\right)(\cos t)^{n-2j}(-1)^j (\sin t) ^{2j}}+i \sum_{j=0}^{\frac{ n-1 }{ 2 }}(\cos t)^{n-(2j+1)}(-1)^j)(\sin t) ^{2j+1}\]

ganeshie8 (ganeshie8):

Yes, but we don't care about the imaginary part here

OpenStudy (richley):

Yes that makes sense, but if we just replace one with the other then we get our initial formula

ganeshie8 (ganeshie8):

\[\color{Red}{\cos(nt)~~~=~~\sum_{j=0}^{\frac{ n-1 }{ 2 }}\left(\begin{matrix}n \\ 2j\end{matrix}\right)(\cos t)^{n-2j}(-1)^j (\sin t) ^{2j}}\]

ganeshie8 (ganeshie8):

Since you want the formula to contain only powers of cosine, you may replace \(\color{red}{(\sin t)^{2j}}\) by \(\color{Red}{(1-\cos^2 t)^j}\) .

OpenStudy (richley):

oh I see

OpenStudy (richley):

Alright I got that down.

OpenStudy (richley):

hello?

OpenStudy (richley):

2) Is now solved, I just need help with #3

OpenStudy (richley):

@ganeshie8

ganeshie8 (ganeshie8):

#3 is pretty straightforward after finishing #2 right ?

OpenStudy (richley):

Using the result of problem 2, express: cos5t in terms of powers of cost (simplest form) sin5t in terms of powers of sint (simplest form)

OpenStudy (richley):

Once I get a start on these problems I can normally finish them. I just never know where to start.

ganeshie8 (ganeshie8):

\[\color{Red}{\cos(nt)~~~=~~\sum_{j=0}^{\frac{ n-1 }{ 2 }}\left(\begin{matrix}n \\ 2j\end{matrix}\right)(\cos t)^{n-2j}(-1)^j (\sin t) ^{2j}}\] replace n by 5

ganeshie8 (ganeshie8):

\[\color{Red}{\cos(5t)~~~=~~\sum_{j=0}^{\frac{ 5-1 }{ 2 }}\left(\begin{matrix}5 \\ 2j\end{matrix}\right)(\cos t)^{5-2j}(-1)^j (\sin t) ^{2j}}\]

ganeshie8 (ganeshie8):

\[\color{Red}{\cos(5t)~~~=~~\sum_{j=0}^{2}\left(\begin{matrix}5 \\ 2j\end{matrix}\right)(\cos t)^{5-2j}(-1)^j (\sin t) ^{2j}}\]

OpenStudy (richley):

Oh, so I replace the n with the 5?

OpenStudy (richley):

And now do I just expand and solve out for the 3 sums, j=0, j=1, and j=2?

ganeshie8 (ganeshie8):

\[\begin{align} &\color{Red}{\cos(5t)} \\~\\ &=\color{red}{\sum_{j=0}^{2}\dbinom{5}{2j}(\cos t)^{5-2j}(-1)^j (\sin t) ^{2j}}\\~\\ &=\color{red}{ \dbinom{5}{0}(\cos t)^{5}(-1)^j (\sin t) ^{0}} + \color{red}{ \dbinom{5}{2}(\cos t)^{3}(-1)^1 (\sin t) ^{2}} + \color{red}{ \dbinom{5}{4}(\cos t)^{1}(-1)^2 (\sin t) ^{4}}\\~\\ \end{align}\]

OpenStudy (richley):

Wait why did it go to \[\left(\begin{matrix}5 \\ 4\end{matrix}\right)\]

OpenStudy (richley):

Oh

OpenStudy (richley):

Nevermind

OpenStudy (richley):

Thats 2j

ganeshie8 (ganeshie8):

You may simplify. Use : \[\dbinom{n}{r} = \dfrac{n!}{(n-r)!*r!}\]

ganeshie8 (ganeshie8):

Also convert sin to cos using sin^2 = 1-cos^2

OpenStudy (richley):

What does "r" represent?

OpenStudy (richley):

oh, the bottom term.